OpenStudy (anonymous):

Use upper and lower sums to approximate the area of the region using the given number of subintervals (of equal width). (Round your answers to three decimal places.) y = sqrt(6x)

OpenStudy (zzr0ck3r):

what interval? what number of intervals?

OpenStudy (anonymous):

0 - 1 4 rectangles

OpenStudy (anonymous):

oh now we can do it

OpenStudy (anonymous):



OpenStudy (anonymous):

each base has length \(.25\) lower endpoints are \(\{0,.25,.5,.75\}\)

OpenStudy (anonymous):

so left hand one will be \[\left(\sqrt{6\times 0}+\sqrt{6\times .25}+\sqrt{6\times .5}+\sqrt{6\times .75}\right)\times .25\]

OpenStudy (anonymous):

and a calculator

OpenStudy (anonymous):

right hand one is similar, but the \(\sqrt{6\times 0}\) will be gone, and replaced by \(\sqrt{6\times 1}\)